Given P(A),find P (not A)
Events A and B are dependent.P(A) = 5/12 and P(B given A) = 6/11. Find (A and B)
Thank You for the straightfoward and correct answers you all send me.

Answer by vleith(2983)

If the probability of A occurring is x%, then the probability of "not A" is 100-x%.
If the probability of A is 'x out of y', the probability of 'not A', is '(y-x) out of y.
The net result is the sum of the two probabilities must be one (or 100%)
Bayes Theorem states:
P(A and B) = P(B)* P(A|B) = P(A)*P(B|A)
You are given P(A) and P(B|A) - you can do the math :-D